# Digital Zero-Noise Extrapolation with Quantum Circuit Unoptimization

### v2

@@ 4 Discussion
We explored using the quantum circuit unoptimization routines introduced in [35] as a noise-scaling method for the zero-noise extrapolation (ZNE) quantum error mitigation technique. To validate our approach, we applied this noise-scaling method to random quantum volume and QAOA circuits, supported by numerical simulations, to demonstrate its effectiveness. These results highlight the potential of quantum circuit unoptimization to introduce controlled noise amplification while preserving the underlying logical functionality of the circuits.
It is worthwhile to consider the trade-offs of quantum circuit unoptimization as a noise-scaling technique compared to other known techniques, such as unitary folding. One of the primary benefits of quantum circuit unoptimization is that, as more iterations of the unoptimization are performed, it becomes progressively more difficult for the original circuit to be recovered via compilers running on real quantum device backends. This property is desirable because it ensures that the noise amplification introduced by the unoptimization process remains intact during execution, preventing compiler optimizations from inadvertently simplifying the circuit. In contrast, unitary folded circuits can sometimes be undone by the compiler, particularly when folding operations are explicitly designed to be reversible and are thus more susceptible to simplification during standard quantum circuit compilation. By comparison, the structural complexity introduced by unoptimization makes it more resistant to such optimizations. Note however that the circuit unoptimization routine could likely be identified and reversed by a compiler that was specifically designed to identify these circuit structures – it appears that current quantum circuit compilers are not able to identify and reverse these structures [35].
While this increased circuit complexity provides advantages in preserving noise scaling, it also introduces potential challenges. Most current quantum computers provide options to disable compiler optimizations, mitigating the risk of unwanted circuit simplifications. However, even when optimizations are turned off, certain low-level adjustments made by the compiler may still partially reverse aspects of the unoptimization process, potentially leading to less predictable noise amplification factors (for example, if the backend removes identity gates that introduce errors due to idling time). This unpredictability arises because unoptimization alters the circuit structure in more challenging ways to model or control than the more straightforward unitary folding approach. The original circuit structure is preserved in unitary folding, and the noise amplification factor is more precisely determined, making it a more predictable method in many scenarios.
Using quantum circuit unoptimization means a slight change in how ZNE is applied. Typically, when using ZNE, the user chooses a noise amplification factor, and then the algorithm attempts to produce a circuit representation with that level of noise amplification. When using quantum circuit unoptimization, instead, the noise scale factor (eq. (3)) is computed based on the circuits generated by the unoptimization, and that value is then used for the extrapolation. Figure 2 shows that the circuit complexity scaling is predictable, which means that a noise amplification factor in principle could be a priori requested. The closest integer required number of circuit unoptimization iterations could then be selected to produce that noise amplification amount. However, this study used a more precise measurement of the noise scale factor post-circuit unoptimization (including post Qiskit transpiler optimization).
Another consideration is the relative complexity of implementation. Unitary folding is conceptually simple and requires minimal additional overhead to apply. By contrast, quantum circuit unoptimization involves multiple steps, including gate insertion, swapping, decomposition, and synthesis, all of which must be implemented carefully to ensure correctness and compatibility with the input circuit. To partially offset this complexity, we provide a fully self-contained implementation of quantum circuit unoptimization in Python, compatible with IBM Qiskit circuits. This implementation can be adapted to other quantum circuit libraries, ensuring accessibility for researchers and practitioners interested in exploring unoptimization as a noise-scaling technique.
@@
An advantage of quantum circuit unoptimization compared to existing noise amplification methods is that an exponentially growing number of potential, distinct, circuit variants can be generated for a particular noise amount. Most other existing methods allow only one or a limited number of circuit variants to be generated. Generating many circuit variants may facilitate better noise averaging across many of these circuit variants on quantum hardware with significantly biased noise.
@@
The primary limitation of quantum circuit unoptimization for noise amplification in ZNE is the highly variable changes to circuit structure. These circuit changes can result in very different error propagation compared to the unmodified circuit. However, the previously mentioned property that many random (equivalent) circuit variants can be generated with quantum circuit unoptimization to mitigate this issue via noise averaging across many instances with different circuit structures. Therefore, in practice on noisy quantum computers, we recommend averaging over many random circuits for each noise scale factor. A possible future subject of study is developing other circuit unoptimization algorithms, in addition to [35], which creates a quantum circuit structure that minimizes highly biased error propagation (e.g., a circuit structure that approximately mimics the original circuit structure).
@@
Finally, it is essential to note that while quantum circuit unoptimization offers certain advantages for specific use cases, there is no theoretical reason to expect it to outperform more conventional approaches to perform consistently noise scaling, such as unitary folding or pulse stretching. Each method has unique strengths and limitations, and the choice of technique should depend on the specific application, the noise characteristics of the hardware, and the desired level of control over circuit structure and complexity.
@@ 5 Software
The software implementing the quantum circuit unoptimization recipe, its application as a noise-scaling technique for ZNE, as well as all data and plots used in this work are available in a public GitHub repository
@@ Appendix A Appendix
The $U$ and $U_{3}$ gates are equivalent in these circuit diagrams due to the legacy naming convention from earlier versions of Qiskit and represents the most general single-qubit rotation. They are parameterized by three angles $\theta$, $\phi$, and $\lambda$, and their matrix representation is given by

### v3

@@ 4 Discussion
We explored using the quantum circuit unoptimization routines introduced in [35] as a noise-scaling method for the zero-noise extrapolation (ZNE) quantum error mitigation technique. To validate our approach, we applied this noise-scaling method to random quantum volume and QAOA circuits, using numerical simulations, to demonstrate its effectiveness. These results show that quantum circuit unoptimization can provide controlled noise amplification while preserving the underlying logical unitary of the digital circuit.
@@
An advantage of quantum circuit unoptimization compared to existing noise amplification methods is that an exponentially growing number of distinct circuit variants can be generated for a particular noise level. All other existing methods allow only one or a small number of circuit variants to be generated for a given noise level. Generating many circuit variants facilitates better noise averaging, which could be specifically useful on quantum computers with biased sources of error.
@@
The primary limitation of quantum circuit unoptimization for noise amplification in ZNE is the highly variable changes to circuit structure. These circuit changes can result in very different error propagation compared to the unmodified circuit. However, the previously mentioned property that many random (equivalent) circuit variants can be generated with quantum circuit unoptimization to mitigate this issue via noise averaging across many instances with different circuit structures. Therefore, in practice on noisy quantum computers, we recommend averaging over many random circuits for each noise scale factor. A possible future subject of study is developing other circuit unoptimization algorithms, in addition to [35], which create a quantum circuit structure that minimizes highly biased error propagation (e.g., a noise-amplified circuit structure that approximately mimics the original circuit structure).
It is worthwhile to consider the trade-offs of quantum circuit unoptimization as a noise-scaling technique compared to other known techniques, such as unitary folding. One of the benefits of quantum circuit unoptimization is that, as more iterations of the unoptimization are performed, it becomes progressively more difficult for the original circuit to be recovered by quantum computer backend compilers. This property is desirable because it ensures that the noise amplification introduced by the unoptimization process remains intact during execution, preventing compiler optimizations from inadvertently simplifying the circuit. In contrast, unitary folded circuits can sometimes be undone by the compiler, particularly when folding operations are explicitly designed to be reversible and are thus more susceptible to simplification during standard quantum circuit compiler optimization. By comparison, the structural complexity introduced by unoptimization makes it more resistant to such optimizations. However, the original circuit structure is preserved in unitary folding, and the noise amplification factor is more precisely determined, making it a more predictable method in many scenarios. Note, however, that the circuit unoptimization routine could likely be identified and reversed by a compiler that was specifically designed to identify these circuit structures – it appears that current quantum circuit compilers are not able to identify and reverse these structures [35]. Most current quantum computers provide options to disable backend compiler optimizations, mitigating the risk of unwanted circuit simplifications. However, even when optimizations are turned off, some low-level adjustments made by the compiler may still partially reverse aspects of noise amplification methods used in ZNE, potentially leading to less predictable noise amplification factors (for example, if the backend removes identity gates that introduce errors due to idling time).
Using quantum circuit unoptimization means a slight change in how ZNE is applied. Typically, when using ZNE, the user chooses a noise amplification factor, and then the algorithm attempts to produce a circuit representation with that level of noise. When using quantum circuit unoptimization, instead, the noise scale factor (eq. (3)) is computed based on the circuits generated by the unoptimization, and that value is then used for the extrapolation. Figure 2 shows that the circuit complexity scaling is predictable, which means that a noise amplification factor in principle could be a priori requested. The closest integer required number of circuit unoptimization iterations could then be selected to produce that noise amplification amount. However, this study used a more precise measurement of the noise scale factor post-circuit unoptimization (including post Qiskit transpiler optimization).
Another consideration is the relative complexity of implementation. Unitary folding is conceptually simple and requires minimal additional overhead to apply. By contrast, quantum circuit unoptimization involves multiple steps, including gate insertion, swapping, decomposition, and synthesis, all of which must be implemented carefully to ensure correctness. To partially offset this complexity, we provide a fully self-contained implementation of quantum circuit unoptimization in Python 3, compatible with IBM Qiskit circuits. This implementation can be adapted to other quantum circuit libraries, ensuring accessibility for researchers and practitioners interested in exploring unoptimization as a noise-scaling technique.
Importantly, note that there is always a core assumption underpinning our method and ZNE in general: the existence of a predictable relationship between the noise-scale factor and the measured expectation value, typically able to be modeled by a low-order polynomial. Our results demonstrate that this assumption holds in the numerical simulations we show, however it may break down for other problem classes, particularly near phase transitions or in algorithms with unstable error propagation. It would be interesting for future work to characterize classes of problems and noise models for which quantum circuit unoptimization ZNE might break down.
@@
Finally, it is essential to note that there is no theoretical reason to expect quantum circuit unoptimization to outperform standard approaches for noise scaling such as unitary folding or pulse stretching. Each method has unique strengths and limitations, and the choice of technique depends on things such as the noise characteristics of the quantum hardware.
@@ 5 Software
The software implementing the quantum circuit unoptimization recipe, its application as a noise-scaling technique for ZNE, as well as all data and plots used in this work is available in a public GitHub repository
@@ Appendix A Appendix
The $U$ and $U_{3}$ gates are equivalent in these circuit diagrams due to the legacy naming convention from earlier versions of Qiskit, and represents the most general single-qubit rotation. They are parameterized by three angles $\theta$, $\phi$, and $\lambda$, and their matrix representation is given by